# M2SVid: End-to-End Inpainting and Refinement for Monocular-to-Stereo Video Conversion

### v1

@@ 6.2 Ablation Study
Full-attention at disoccluded pixels (Sec. 5.3):  While this contribution doesn’t impact metrics in Tab. 3, in Fig. 5, we can observe that computing full-attention in the inpainting regions prevents hallucination of thin structures.
End-to-end training (Sec. 5.4):  In Tab. 4, we ablate our training strategy. When comparing the model trained with the standard diffusion loss (Eq. 1) with 25 (A) or 1 (B) diffusion step at inference, the single step model leads to more blurry results than (A), as evidenced by the 12% and 20% improvement in LPIPS on Stereo4D and Ego4D respectively. End-to-end training with only latent space supervision (Eq. 2) (C) only slightly improves the LPIPS metric. Our final training loss (Eq. 5) (D), including LPIPS and L1 losses in image space, largely improves the image sharpness, as evidenced by the 24% and 22% improvement in LPIPS on Stereo4D and Ego4D compared to (C). Notably, we outperform the standard diffusion loss with 25 inference steps (A) on all metrics and datasets.
@@ 7 Conclusion
We introduce an end-to-end approach for stereoscopic video inpainting and refinement in this work. We utilize Stable Video Diffusion as our base model, and extend it to input the warped right view, discocclusion mask, and the input left view as conditioning. Next, we modify the attention layers in SVD to compute full attention for the discoccluded pixels in order to improve inpainting quality. Crucially, we perform the video refinement using a single denoising step, enabling end-to-end training with image space losses. Qualitative as well as quantitative experiments demonstrate that our method clearly outperforms prior state-of-the-art methods for the monocular-to-stereo video conversion task.
@@
The supplementary material provides more details about the datasets, implementation, and results. Appendix A provides a brief background on Diffusion Models. More details about our training and evaluation datasets are provided in Appendix B. Appendix C provides further implementation details, while Appendix D contains more detailed results and analysis. Finally, Appendix E provides details on inference for high-resolution long videos. We also include the results generated by our methods as well as the compared approaches, on the 21 videos used for our user study. You can view each of them independently, or open the index.html file in a browser. We also provide additional results generated by our model on high-resolution long videos as described in Appendix E.
@@ Appendix A Background on Diffusion Models
DDPMs:  Denoising Diffusion Probabilistic Models (DDPMs) [20] are generative models trained to map a simple noise distribution $p_{T}$ to the data distribution $p_{0}$, by reversing a stochastic forward process $p_{t}$, $t=1,...,T$. This forward process gradually adds small amounts of Gaussian noise with variance $\beta_{t}$, ensuring the reverse process can be approximated as a Gaussian distribution. The forward process might be also defined by $\boldsymbol{\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\mathbf{\epsilon}}$, where $\boldsymbol{\mathbf{x}}_{0}$ is a data sample, $\boldsymbol{\mathbf{\epsilon}}\sim\mathcal{N}(\text{0},\text{I})$ is a noise, and $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{\tau=1}^{t}\alpha_{\tau}$. A denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t)$ is trained to progressively remove noise from the input, by predicting $\boldsymbol{\mathbf{x}}_{t-1}$ from $\boldsymbol{\mathbf{x}}_{t}$ and timestamp $t$, effectively reversing the diffusion.
DDIMs:  Denoising Diffusion Implicit Models (DDIMs) [70] extend DDPMs by introducing non-Markovian diffusion processes, preserving the original training objective while enabling much faster inference with as few as 25 or 50 steps.
LDMs:  To facilitate training, Latent Diffusion Models (LDMs) [60] operate in the latent space of a Variational Autoencoder (VAE) [34], which consists of an encoder $E$ and a decoder $D$. The VAE is trained separately and remains frozen during diffusion model training. By working in the latent space, where $D(E(x))\approx x$, LDMs significantly reduce computational complexity while preserving essential data structure. The diffusion model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}$ is then trained directly in this lower-dimensional latent space.
Conditional diffusion models:  Finally, conditional diffusion models [62, 88] extend diffusion models by conditioning the denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t,c)$ on additional input $c$, such as text [62], images [62], or pose and depth maps [88].
@@ B.2 Ego4D
Ego4d Pre-processing:  We perform rectification for each long video as described in Sec. 5.5 of the main paper. For video rectification, we uniformly sample 200 frames from each video. Then, we split the videos into 5-second subsequent video clips and compute shifts and disparity at the clip level. Finally, to ensure high-quality rectified data, we perform LoFTR feature matching [71] again and filter point pairs by computing the fundamental matrix with RANSAC [13], applied to frames sampled at 1 FPS from each clip. We then filter out clips where the matched points exhibit more than 2 pixels of vertical disparity.
@@ B.3 DAVIS
DAVIS: The dataset consists of 50 videos, each up to 6 seconds long, with a resolution of $1024\times 1920$. The videos are highly dynamic, featuring moving cameras and one or more moving objects (e.g., BMX riding, breakdancing, etc.). Note that the dataset only contains monocular videos, without an ground truth depth or right view. Hence we use this dataset solely for qualitative comparisons and user studies.
Figure 8: The users were shown the above intructions during the study.
@@
SVG [10] - (training free) 50 steps 25.6 0.926 0.217 38.1 0.994 0.014 26.3 0.940 0.190
StereoCrafter [68] private dataset 25 steps 24.9 0.909 0.242 38.2 0.995 0.012 25.5 0.922 0.217
StereoCrafter [68] private dataset 1 step 25.3 0.911 0.262 38.9 0.996 0.014 25.8 0.923 0.234
@@ C.1 Conditioning
Model training:  We initialize our M2SVid model using the Stable Video Diffusion model [4]. We train the model for 300k iterations with a batch size of 16 and a learning rate of $2\times 10^{-6}$. Due to GPU memory limitations, during training, we sample batches with frame number and resolutions of $4\times 512\times 512$, $16\times 256\times 256$, and $25\times 192\times 192$, utilize gradient checkpointing[6] and train the model in float16 precision. We employ random resized cropping, using a resize scale in the range of [0.3, 1.0], for data augmentation. We preprocess the disocclusion masks using the closing morphological transformation to fill small holes inside the inpainting regions with a kernel size of 11 during training and testing.
Depth estimation with DepthCrafter:  During inference for depth prediction, we utilize the DepthCrafter [22] method, the state-of-the-art in diffusion-based video-depth estimation approach. This model can accommodate up to 110 frames and out temporally consistent depth. However, any depth model might be used at this step, for example, efficient 1-step depth diffusion models as in [16]. We predict depth in chunks of 110 frames with an overlap of 25 frames. Depth was predicted on frames resized to a resolution of 1024, followed by upscaling to the original size.
@@ D.1 State-of-the-art Methods
SVG is a training-free method that utilizes a frozen diffusion model to inpaint regions within the mask while preserving the other regions. To achieve spatial-temporal consistency, the method employs inpainting of 8 uniformly sampled views between the given left view and the required right view at the same time, increasing the computational burden. We use 50 denoising steps as recommended. SVG [10] has been compared to state-of-the-art video inpainting [93, 38] and dynamic novel view synthesis [45, 39] methods, significantly outperforming them in temporal consistency and overall inpainting quality. This highlights the fundamental differences between stereo video inpainting, classical video inpainting, and novel view synthesis, revealing a substantial gap between these tasks. Due to this disparity, we exclude classical video inpainting and novel view synthesis methods from our comparisons.
StereoCrafter [68] is a concurrent method and the closest baseline to our approach. It fine-tunes a diffusion model, SVD, to perform inpainting given a warped view and a mask in a diffusion-based denoising setup, trained on a privately collected internet dataset. Since the method does not specify the number of denoising steps, we use 25 steps, as is commonly done [22].
@@
D.2 User Study
In order to perform our human perception user study to rank the different methods, we selected a set of 21 videos. We tried to ensure a diverse collection of videos covering different types of subjects, video quality, and motion pattern. We avoided including without with very large motions since we found it very hard to compare the methods in such cases.
During the study, the participants were provided HTML pages containing the links to the generated predictions, as shown in Figure 7. Furthermore, the users were shown detailed instructions about the study at the top of the HTML page. The exact instructions are shown in Figure 8. In particular, the users were provided a brief background about the problem statement and the method. They were asked to ranking the quality of the different methods based on temporal consistency, image quality (i.e. sharpness), lack of artifacts, etc. In order to miminize noise, the users were allowed to assign same ranks to two or more methods in case they were indistinguishable.
D.3 State-of-the-art Comparison
For quantitative evaluation, we use PSNR, MS-SSIM, and LPIPS [89] metrics, computed independently for each video and averaged over the full dataset. Additionally, we report these metrics separately for the disoccluded and non-disoccluded regions in Tab. 5. To achieve this, we mask all pixels outside the considered region with white pixels, compute the metrics at the video level, and then average them over the dataset.
@@ D.4 Ablations
In Tabs. 6 and 7, we further report metrics for the disoccluded regions for our ablation studies. We find that incorporating the left view as an additional conditioning signal (Tab. 6) results in a slight improvement in inpainting performance across both datasets. Additionally, end-to-end training with latent space supervision (Tab. 7) achieves nearly the same inpainting performance as a model trained with the standard diffusion paradigm and evaluated with 1-step inference. However, end-to-end training with image-based loss leads to a substantial drop in LPIPS, particularly on the Ego4D dataset.
@@ Appendix E Inference on Arbitrary Length and Resolution
To achieve stereo video conversion for videos of arbitrary length, we fine-tune our M2SVid model for 20K iterations using the auto-regressive modeling approach proposed in [68]. Specifically, during training, we replace the first $n$ frames of $V_{r}^{warp}$ with ground truth frames from the right video $V_{r}$, where $n$ is randomly sampled between 00 and $N$. During inference on long videos, we use the last $m$ frames generated in the current round as input for the next round, ensuring seamless stitching between consecutive rounds. For inference with arbitrary resolution, following [68], we utilize tiled diffusion. Specifically, we divide high-resolution videos into overlapping blocks and perform stereo video conversion on each block independently. The overlapping regions are blended in the latent space before VAE decoding to ensure seamless transitions between blocks. We provide examples generated by our model on high-resolution long videos in the supplementary material zip file.

### v2

@@ 6.2 Ablation Study
Full-attention at disoccluded pixels (Sec. 5.3):  While this contribution doesn’t impact metrics in Tab. 4 (likely due to the limited number of complex scenes in the test set), our qualitative analysis shows benefits for dynamic scenes with camera motion, where both foreground and background move. In Fig. 6, we observe that our method prevents hallucinations (top example) and enables correct inpainting (bottom) of thin structures. More results are in Appendix.
End-to-end training (Sec. 5.4):  In Tab. 5, we ablate our proposed end-to-end training strategy. When comparing the model trained with the standard diffusion loss (Eq. 1) with 25 (A) or 1 (B) diffusion step at inference, the single step model leads to more blurry results than (A), as evidenced by the 12% and 20% decrease in LPIPS on Stereo4D and Ego4D respectively. End-to-end training with only latent space supervision (Eq. 2) (C) only slightly improves the LPIPS metric. Our final training loss (Eq. 5) (D), including LPIPS and L1 losses in image space, largely improves the image sharpness, as evidenced by the 24% and 22% improvement in LPIPS on Stereo4D and Ego4D compared to (C). Notably, we outperform the standard diffusion loss with 25 inference steps (A) on all metrics and datasets.
@@ 7 Conclusion
We introduce an end-to-end approach for stereoscopic video inpainting and refinement in this work. First, we extend the SVD model to take the input left video, warped right video, and disocclusion mask as conditioning. Next, we modify the attention layers in SVD to compute full attention for the discoccluded pixels in order to improve inpainting quality. Crucially, we perform the video refinement using a single denoising step, enabling end-to-end training with image space losses. Qualitative and quantitative experiments, as well as user studies, demonstrate that our method clearly outperforms prior state-of-the-art methods for the monocular-to-stereo video conversion task.
@@
The supplementary material provides more details about the datasets, implementation, and results. Appendix A provides a brief background on Diffusion Models. More details about our training and evaluation datasets are provided in Appendix B. Appendix C provides further implementation details as well as additional information on user studies, while Appendix D contains more detailed results and analysis. Appendix E provides details on inference for high-resolution long videos. Finally, Appendix F discusses the limitations of our approach.
@@ Appendix A Background on Diffusion Models
DDPMs:  Denoising Diffusion Probabilistic Models (DDPMs) [ho2020denoising] are generative models trained to map a simple noise distribution $p_{T}$ to the data distribution $p_{0}$, by reversing a stochastic forward process $p_{t}$, t=1,,Tt=1,...,T. This forward process gradually adds small amounts of Gaussian noise with variance $\beta_{t}$, ensuring the reverse process can be approximated as a Gaussian distribution. The forward process might be also defined by $\boldsymbol{\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\mathbf{\epsilon}}$, where $\boldsymbol{\mathbf{x}}_{0}$ is a data sample, $\boldsymbol{\mathbf{\epsilon}}\sim\mathcal{N}(\text{0},\text{I})$ is a noise, and $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{\tau=1}^{t}\alpha_{\tau}$. A denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t)$ is trained to progressively remove noise from the input, by predicting $\boldsymbol{\mathbf{x}}_{t-1}$ from $\boldsymbol{\mathbf{x}}_{t}$ and timestamp $t$, effectively reversing the diffusion.
DDIMs:  Denoising Diffusion Implicit Models (DDIMs) [song2020denoising] extend DDPMs by introducing non-Markovian diffusion processes, preserving the original training objective while enabling much faster inference with as few as 25 or 50 steps.
LDMs:  To facilitate training, Latent Diffusion Models (LDMs) [rombach2022high] operate in the latent space of a Variational Autoencoder (VAE) [kingma2013auto], which consists of an encoder $E$ and a decoder $D$. The VAE is trained separately and remains frozen during diffusion model training. By working in the latent space, where $D(E(x))\approx x$, LDMs significantly reduce computational complexity while preserving essential data structure. The diffusion model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}$ is then trained directly in this lower-dimensional latent space.
Conditional diffusion models:  Finally, conditional diffusion models [saharia2022photorealistic, zhang2023adding] extend diffusion models by conditioning the denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t,c)$ on additional input $c$, such as text [saharia2022photorealistic], images [saharia2022photorealistic], or pose and depth maps [zhang2023adding].
Figure 8: The users were shown the above intructions during the study.
@@ B.2 Ego4D
Ego4d Pre-processing:  We perform rectification for each long video as described in Sec. 5.5 of the main paper. For video rectification, we uniformly sample 200 frames from each video. After rectifying the videos, we spatially crop the video to discard boundary regions with missing pixels. Then, we split the videos into 5-second subsequent video clips and compute shifts and disparity at the clip level. Finally, to ensure high-quality rectified data, we perform LoFTR feature matching [sun2021loftr] again and filter point pairs by computing the fundamental matrix with RANSAC [fischler1981random], applied to frames sampled at 1 FPS from each clip. We then filter out clips where the matched points exhibit more than 2 pixels of vertical disparity.
@@ B.3 DAVIS
DAVIS: The dataset consists of 50 videos, each up to 6 seconds long, with a resolution of $1024\times 1920$. The videos are highly dynamic, featuring moving cameras and one or more moving objects (e.g., BMX riding, breakdancing, etc.). Note that the dataset only contains monocular videos, without a ground truth depth or ground truth right view. Hence we use this dataset solely for qualitative comparisons and user studies.
@@
SVG [dai2024svg] - (training free) 50 steps 25.6 0.926 0.217 38.1 0.994 0.014 26.3 0.940 0.190
StereoCrafter [shi2024stereocrafter] private dataset 25 steps 24.9 0.909 0.242 38.2 0.995 0.012 25.5 0.922 0.217
StereoCrafter [shi2024stereocrafter] private dataset 1 step 25.3 0.911 0.262 38.9 0.996 0.014 25.8 0.923 0.234
@@ C.1 Conditioning
Model training:  We initialize our M2SVid model using the Stable Video Diffusion model (stable-video-diffusion-img2vid-xt version) [blattmann2023stable]. We train the model for 300k iterations with a batch size of 16 and a learning rate of $2\times 10^{-6}$. Due to GPU memory limitations, during training, we sample batches with frame number and resolutions of $4\times 512\times 512$, $16\times 256\times 256$, and $25\times 192\times 192$, utilize gradient checkpointing[chen2016training] and train the model in float16 precision. We employ random resized cropping, using a resize scale in the range of [0.3, 1.0], for data augmentation. We preprocess the disocclusion masks using the closing morphological transformation to fill small holes inside the inpainting regions with a kernel size of 11 during training and testing.
Depth estimation with DepthCrafter:  During inference for depth prediction, we utilize the DepthCrafter [hu2024depthcrafter] method, the state-of-the-art in diffusion-based video-depth estimation approach. This model can accommodate up to 110 frames and output temporally consistent depth. However, any depth model might be used at this step, for example, efficient 1-step depth diffusion models as in [garcia2024fine]. We predict depth in chunks of 110 frames with an overlap of 25 frames. Depth was predicted on frames resized to a resolution of 1024, followed by upscaling to the original size.
C.2 Desktop User Study
In order to perform our desktop human perception user study to rank the different methods, we selected a set of 21 videos. We tried to ensure a diverse collection of videos covering different types of subjects, video quality, and motion pattern. We avoided including without with very large motions since we found it very hard to compare the methods in such cases.
During the study, the participants were provided HTML pages containing the links to the generated predictions, as shown in Figure 7. Furthermore, the users were shown detailed instructions about the study at the top of the HTML page. The exact instructions are shown in Figure 8. In particular, the users were provided a brief background about the problem statement and the method. They were asked to rank the quality of the different methods based on temporal consistency, image quality (i.e. sharpness), lack of artifacts, etc. In order to miminize noise, the users were allowed to assign same ranks to two or more methods in case they were indistinguishable.
C.3 VR Headset User Study
For the VR headset user study, the participants were shown the stereoscopic videos generated by our method and StereoCrafter (25 denoising steps) on a video player that supports side-by-side (3D) content. We used the same 21 videos that were used in the Desktop user study. The participants were then asked to rank the methods based on general viewing experience, while allowing for tied ratings. Note that all the videos were anonymized and the order of showing the methods was also randomized. We had 5 distinct participants in the user study (3 male and 2 female), each of whom provided rankings for all 21 videos.
@@ D.1 State-of-the-art Methods
SVG is a training-free method that utilizes a frozen diffusion model to inpaint regions within the mask while preserving the other regions. To achieve spatial-temporal consistency, the method employs inpainting of 8 uniformly sampled views between the given left view and the required right view at the same time, increasing the computational burden. We use 50 denoising steps as recommended. SVG [dai2024svg] has been compared to state-of-the-art video inpainting [zhou2023propainter, li2022towards] and dynamic novel view synthesis [liu2023robust, li2023dynibar] methods, significantly outperforming them in temporal consistency and overall inpainting quality. This highlights the fundamental differences between stereo video inpainting, classical video inpainting, and novel view synthesis, revealing a substantial gap between these tasks. Due to this disparity, we exclude classical video inpainting and novel view synthesis methods from our comparisons.
StereoCrafter [shi2024stereocrafter] is a concurrent method and the closest baseline to our approach. It fine-tunes a diffusion model, SVD, to perform inpainting given a warped view and a mask in a diffusion-based denoising setup, trained on a privately collected internet dataset. Since the method does not specify the number of denoising steps, we use 25 steps, as is commonly done [hu2024depthcrafter].
@@
D.2 State-of-the-art Comparison
For quantitative evaluation, we use PSNR, MS-SSIM, and LPIPS [zhang2018perceptual] metrics, computed independently for each video and averaged over the full dataset. Additionally, we report these metrics separately for the disoccluded and non-disoccluded regions in Tab. 6. To achieve this, we mask all pixels outside the considered region with white pixels, compute the metrics at the video level, and then average them over the dataset.
@@ D.2 State-of-the-art Comparison
D.3 Full-attention at Disoccluded Pixels.
In Figs. 10, 11 and 12, we provide further qualitative evaluation of our proposed full attention at disoccluded pixels. We find that full attention is particularly beneficial for scenes with complex backgrounds and strong camera movement, where both foreground and background pixels move and correct inpainting requires the model to “copy” disoccluded pixels from different spatial locations in other frames. For example, in Fig. 10, the handrail of the stairs in frame 4 can only be correctly inpainted by using information from a different spatial location in frame 9 due to camera movement. In Fig. 11, due to the strong movement of the motorcyclist behind the central rider, to inpaint the region near the helmet in frame 11, the model has to use data from a strongly shifted spatial location in frame 10. Finally, in Fig. 12, while standard factorized attention hallucinated the reins in frame 16 at the same spatial location as in the previous frame, creating a visual artifact as if the reins split into two pieces, full attention at disoccluded pixels prevents this hallucination and correctly inpaints the region.
@@ D.4 Ablations
In Tabs. 7 and 8, we further report metrics for the disoccluded regions for our ablation studies. We find that incorporating the left view as an additional conditioning signal (Tab. 7) results in a slight improvement in inpainting performance across both datasets. Additionally, end-to-end training with latent space supervision (Tab. 8) achieves nearly the same inpainting performance as a model trained with the standard diffusion paradigm and evaluated with 1-step inference. However, end-to-end training with image-based loss leads to a substantial drop in LPIPS, particularly on the Ego4D dataset.
@@ D.5 Run-time
[…]
regime training steps Stereo4D Ego4d
Sampling $t$ $L_{latent}$ 1 ($t=T$) 0.242 0.320
5 0.217 0.278
25 0.215 0.267
Fixed $t=T$ $L_{latent}+L_{\text{LPIPS}}+L_{\text{L1}}$ 1 ($t=T$) 0.179 0.244
5 0.200 0.260
25 0.243 0.298
Table 10: The effect of the number of inference steps on model performance. In standard training with sampled $t$, more inference steps improve results, while 1-step inference performs poorly. In contrast, our model is trained with fixed $t=T$, enabling end-to-end supervision in image space (which is not possible with sampled $t$). This leads to superior 1-step performance, while multi-step inference degrades results due to mismatch with training.
(a) Crop from input left view
(b) VAE encoded-decoded left view
(c) Generated right (stitching area)
(d) Generated right (not stitching area)
Figure 9: Blurriness due to usage of VAE and spatial stitching in high-resolution videos.
Figure 10: Standard factorized attention vs full-attention at dis-occluded pixels (ours). Full attention at dis-occluded pixels helps the model better exploit information from other frames, in particular by correctly inpainting the handrail of the stairs using data from a different spatial location in another frame.
Figure 11: Standard factorized attention vs full-attention at dis-occluded pixels (ours). Full attention at dis-occluded pixels helps the model better exploit information from other frames, in particular by correctly inpainting the region near the helmet, using data from a different spatial location in another frame.
Figure 12: Standard factorized attention vs full-attention at dis-occluded pixels (ours). Full attention at dis-occluded pixels helps the model better exploit information from other frames. In particular, while standard factorized attention hallucinated the reins at the same spatial location as in the previous frame, creating a visual artifact as if the reins split into two pieces, full attention at dis-occluded pixels prevents this hallucination and correctly inpaints the region.
@@
D.6 Number of Denoising Steps
In Tab. 10, we analyze the effect of the number of inference steps on model performance. In a standard training regime, where a timestamp $t$ is randomly sampled during training, one-step inference results in poor performance, while increasing the number of inference steps improves it. In contrast, we explicitly train a one-step model by fixing the timestamp to $t=T$ during training. This enables end-to-end training with an image-space loss, as the U-Net directly predicts denoised latents that can be decoded with a VAE and supervised in image space. Note that training with an image-space loss is not possible in the standard regime with sampled $t$, since when $t\neq T$, the U-Net outputs a mixture of latents and noise (not clean latents), which therefore cannot be decoded with the VAE and supervised in image space. As a result of end-to-end training with a single step, our model achieves the best performance with one-step inference. However, multi-step inference degrades performance, as the model was not trained with $t\neq T$.
@@ Appendix E Inference on Arbitrary Length and Resolution
To achieve stereo video conversion for videos of arbitrary length, we fine-tune our M2SVid model for 20K iterations using the auto-regressive modeling approach proposed in [shi2024stereocrafter]. Specifically, during training, we replace the first $n$ frames of $V_{r}^{warp}$ with ground truth frames from the right video $V_{r}$, where $n$ is randomly sampled between 0 and $N$. During inference on long videos, we use the last $m$ frames generated in the current round as input for the next round, ensuring seamless stitching between consecutive rounds. For inference with arbitrary resolution, following [shi2024stereocrafter], we utilize tiled diffusion. Specifically, we divide high-resolution videos into overlapping blocks and perform stereo video conversion on each block independently. The overlapping regions are blended in the latent space before VAE decoding to ensure seamless transitions between blocks. We provide examples generated by our model on high-resolution long videos in the supplementary material zip file.
Appendix F Limitations
We build our model upon the Stable Video Diffusion model, which allows us to leverage learned video priors obtained through large-scale generative pretraining. Due to the high computational cost of processing video data, SVD employs a VAE to compress videos into a lower-resolution latent space, where denoising is performed. Following this architecture, we first encode the video using a VAE, then apply a U-Net, and finally decode the output back to video space using the same VAE. However, VAE compression may lead to loss of high-frequency details. As illustrated in Fig. 9, even encoding and decoding the left-view videos with the VAE alone can introduce visible blurriness (e.g., the shoe in Figs. 9(a) and 9(b)) and loss of fine details. To support stereo conversion for higher-resolution videos, we additionally apply temporal and spatial stitching, as described in Appendix E. While simple and efficient, stitching may introduce further blurring artifacts, as shown in Figs. 9(c) and 9(d). We plan to address these limitations in future research.